# Fractional binomial distributions induced by the generalized binomial theorem and their applications

### v1

@@ 4.1. The fractional Bernstein operator and its approximating property
Let $\alpha>0$, $n\in\mathbb{N}$, $x\in[0,1]$, and $S_{\alpha,x}^{(n)}$ be a random variable whose probability law is $\mu_{\alpha,x}^{(n)}$. Then, $B_{\alpha,n}f$, $f\in C([0,1])$ can be expressed as
@@
$\hskip 28.45274pt+\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right]$
@@
4.2. The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
It is natural to consider the behaviors of the $k$-fold iterates $B_{\alpha,n}^{k}$ of the $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ as $k\to\infty$. In this section, we prove the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$ based on an argument similar to that in the proof of [Rus04, Theorem 1].
@@ 4.2. The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
which means that $B_{\alpha,n}$ is contractive on $C_{a,b}([0,1])$ for $a,b\in\mathbb{R}$. Then, the contraction principle allows us to find the unique function $f_{\alpha,n}\in C([0,1])$ satisfying $B_{\alpha,n}f_{\alpha,n}=f_{\alpha,n}$, $f_{\alpha,n}(0)=f(0)$, $f_{\alpha,n}(1)=f(1)$, and (4.2). The explicit representation of $f_{1,n}$ was obtained in Proposition 1.6. ∎
4.3. Limit theorems for iterates of the fractional Bernstein operator
@@
Lemma 4.3 (cf. [EK86, Theorem 8.2.8] and [AC94, Theorem 6.2.6]).
@@
We prove that the discrete semigroup defined by the iterates of $B_{\alpha,n}$ converges to $(T_{\alpha}(t))_{t\geq 0}$ as follows.
@@
$\hskip 28.45274pt+n\mathbb{E}\Bigg{[}\int_{0}^{1}(1-t)\left\{f^{\prime\prime}\left(x+t\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)\right)-f^{\prime\prime}(x)\right\}\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\,\mathrm{d}t\Bigg{]}$
$\hskip 28.45274pt+\frac{n}{2Z_{\alpha,x}^{(n)}}\left\{\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}-\frac{1}{\alpha n}x(1-x)\right\}f^{\prime\prime}(x)$
$\hskip 28.45274pt+\frac{1}{2\alpha}\left(\frac{1}{Z_{\alpha,x}^{(n)}}-1\right)x(1-x)f^{\prime\prime}(x)$
$=:\mathrm{I}_{\alpha}^{(n)}(x)+\mathrm{II}_{\alpha}^{(n)}(x)+\mathrm{III}_{\alpha}^{(n)}(x)+\mathrm{IV}_{\alpha}^{(n)}(x).$ (4.5)
@@
$\hskip 28.45274pt+n\|f^{\prime}\|_{\infty}\max_{x\in[r,1-r]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|$
@@
$\hskip 28.45274pt+n\|f^{\prime}\|_{\infty}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)O(e^{-\delta n})$
@@
$\left|\mathrm{II}_{\alpha}^{(n)}(x)\right|$
@@
$\hskip 14.22636pt+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{4}\right|$
@@
$\left|\mathrm{III}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{n}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\|f^{\prime\prime}\|_{\infty}O\left(\frac{1}{n^{2}}\right)=O\left(\frac{1}{n}\right)$ (4.8)
as $n\to\infty$. On the other hand, for $\varepsilon>0$, there exists some sufficiently small $u=u(\varepsilon)>0$ satisfying $x(1-x)<\varepsilon$ for $x\in[0,u]\cup[1-u,1]$. Then, we have
$\left|\mathrm{IV}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\max_{x\in[0,u]\cup[1-u,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\max_{x\in[u,1-u]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
$\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\sup_{\alpha\in K}\max_{x\in[u,1-u]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
@@
as $n\to\infty$, by virtue of (2.13) and Theorem 2.9. Therefore, by combining (4.5) with (4.3), (4.7), (4.8), and (4.9), we obtain
@@
The function $g_{\delta}$ (resp. $\widehat{g}_{\delta}$) is non-negative and continuous on $[0,\varepsilon]$ (resp. $[1-\varepsilon,1]$) with $g_{\delta}(\varepsilon)=0$ (resp. $\widehat{g}_{\delta}(1-\varepsilon)=0$). We also define
@@
$G(0+)$ =+=$\infty$,= + ∞ , $\qquad G(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(0)|,$
$\widehat{G}(0+)$ =+=$\infty$,= + ∞ , $\qquad\widehat{G}(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(1)|$
due to (4.11). Hence, the mean value theorem implies the existences of $\delta_{0},\delta_{1}\in(0,\varepsilon)$ such that $G(\delta_{0})=|f^{\prime}(0)|$ and $\widehat{G}(\delta_{1})=|f^{\prime}(1)|$. We now define a family of continuous functions $(f_{\varepsilon})_{\varepsilon>0}\subset C([0,1])$ by
$f_{\varepsilon}(x):=\begin{cases}f(x)+\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle\int_{x}^{\varepsilon}\int_{y}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s\,\mathrm{d}y&\text{if }x\in[0,\varepsilon),\\ f(x)&\text{if }x\in[\varepsilon,1-\varepsilon],\\ f(x)+\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\int_{1-\varepsilon}^{x}\int_{1-\varepsilon}^{y}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s\,\mathrm{d}y&\text{if }x\in(1-\varepsilon,1].\\ \end{cases}$
@@
$f_{\varepsilon}^{\prime}(x):=\begin{cases}f^{\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle\int_{x}^{\varepsilon}g_{\delta_{0}}(s)\,\mathrm{d}s&\text{if }x\in[0,\varepsilon),\\ f^{\prime}(x)&\text{if }x\in(\varepsilon,1-\varepsilon),\\ f^{\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\int_{1-\varepsilon}^{x}\widehat{g}_{\delta_{1}}(s)\,\mathrm{d}s&\text{if }x\in(1-\varepsilon,1].\\ \end{cases}$
@@
$f_{\varepsilon}^{\prime\prime}(x):=\begin{cases}f^{\prime\prime}(x)+\{\mathrm{sgn}\,f^{\prime}(0)\}\displaystyle g_{\delta_{0}}(x)&\text{if }x\in[0,\varepsilon),\\ f^{\prime\prime}(x)&\text{if }x\in(\varepsilon,1-\varepsilon),\\ f^{\prime\prime}(x)-\{\mathrm{sgn}\,f^{\prime}(1)\}\displaystyle\widehat{g}_{\delta_{1}}(x)&\text{if }x\in(1-\varepsilon,1]\\ \end{cases}$
@@
It follows from Lemma 4.6 and [EK86, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$. Then, by combining this fact with Lemma 4.5, we can apply Trotter’s approximation theorem (cf. [Tro58]) to conclude (4.3). ∎
@@
Because (4.14) has a unique strong solution, $(X_{\alpha,x}(t))_{t\geq 0}$ is the unique solution to the martingale problem for $(\mathcal{L}_{\alpha},\delta_{x})$ for every $x\in[0,1]$. Moreover, we note that the diffusion coefficient $(2\alpha)^{-1}x(1-x)$ of $\mathcal{L}_{\alpha}$ is continuous and bounded on $[0,1]$. Then, we can apply [Kal21, Theorem 32.11] to conclude that
@@
is a Feller semigroup on $C([0,1])$ and $\mathcal{L}_{\alpha}$ uniquely extends to the associated generator. Because a Feller semigroup is uniquely determined by its generator (cf. [Kal21, Lemma 17.5]), we have $T_{\alpha}(t)=\widetilde{T}_{\alpha}(t)$ for all $t\geq 0$. ∎
4.4. The fractional Bernstein operator with $\alpha=\alpha(x)$ being a function of $x$
@@
$\mu_{\alpha(x),x}^{(n)}(\{j\})=\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)},\qquad x\in[0,1],\,j=0,1,2,\dots,n$
@@
$B_{\alpha(\cdot),n}f(x)$ $=\sum_{j=0}^{n}\mu_{\alpha(x),x}^{(n)}(\{j\})f\left(\frac{j}{n}\right)$
@@
for $f\in C^{2}([0,1])$ and $x\in[0,1]$. Then, the operator $(\mathcal{L}_{\alpha(\cdot)},C^{2}([0,1]))$ is also closable on $C([0,1])$, and its closure $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$ generates the Feller semigroup $(T_{\alpha(\cdot)}(t))_{t\geq 0}$ on $C([0,1])$, which is no longer the Wright–Fisher diffusion semigroup in general (see [AC94, Theorems 1.6.11 and 6.2.6]).
By following the proof of Theorem 4.4, we also obtain that the discrete semigroup defined by the iterates of $B_{\alpha(\cdot),n}$ converges to $(T_{\alpha(\cdot)}(t))_{t\geq 0}$.
@@
as $\varepsilon\searrow 0$, which implies that $\mathcal{D}_{0}$ is also a core for $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$. Therefore, it follows from [EK86, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha(\cdot)}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$. Then, Trotter’s approximation theorem allows us to obtain (4.16). ∎
@@
For $x\in[0,1]$, let $(X_{\alpha(\cdot),x}(t))_{t\geq 0}$ be the one-dimensional diffusion process that solves the stochastic differential equation
@@
Here, we note that (4.17) has the unique strong solution by applying the Yamada–Watanabe theorem (cf. [YW71, Theorem 1]). We omit the proof of Proposition 4.9 because it is similar to that of Proposition 4.7.
Acknowledgements. The first-named author was supported by JSPS KAKENHI Grant Number 19H00643. The second-named author was supported by JSPS KAKENHI Grant Number 23K12986.

### v3

@@ 4.1 The fractional Bernstein operator and its approximating property
Let $\alpha>0$, $n\in\mathbb{N}$, $x\in[0,1]$, and $S_{\alpha,x}^{(n)}$ be a random variable whose law is $\mu_{\alpha,x}^{(n)}$. Then, $B_{\alpha,n}f$, $f\in C([0,1])$ can be expressed as
@@
$\quad+\mathbb{E}\left[\left|f\left(\frac{1}{n}S_{\alpha,x}^{(n)}\right)-f(x)\right|\,:\,\left|\frac{1}{n}S_{\alpha,x}^{(n)}-x\right|\geq\rho\right]$
@@
4.2 The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
It is natural to consider the behaviors of the $k$-fold iterates $B_{\alpha,n}^{k}$ of the $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ as $k\to\infty$. In this section, we prove the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$ based on an argument similar to that in the proof of [28, Theorem 1].
@@ 4.2 The Kelisky–Rivlin-type theorem for the fractional Bernstein operator
which means that $B_{\alpha,n}$ is contractive on $C_{a,b}([0,1])$ for $a,b\in\mathbb{R}$. Then, the contraction principle allows us to find the unique function $f_{\alpha,n}\in C([0,1])$ satisfying $B_{\alpha,n}f_{\alpha,n}=f_{\alpha,n}$, $f_{\alpha,n}(0)=f(0)$, $f_{\alpha,n}(1)=f(1)$, and (4.2). The explicit representation of $f_{1,n}$ is obvious from [16, Theorem 1]. ∎
4.3 Limit theorems for iterates of the fractional Bernstein operator
@@
Lemma 4.3 (cf. [7, Theorem 8.2.8] and [2, Theorem 6.2.6]).
@@
We prove that the appropriately scaled operators $(B_{\alpha,n}^{\lfloor nt\rfloor})_{n\in\mathbb{N}_{0}}$, $t\geq 0$, defined by the iterates of $B_{\alpha,n}$ converges to $(T_{\alpha}(t))_{t\geq 0}$ as follows.
@@
$\quad+n\mathbb{E}\Bigg[\int_{0}^{1}(1-t)\left\{f^{\prime\prime}\left(x+t\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)\right)-f^{\prime\prime}(x)\right\}\left(\frac{1}{n}S_{\alpha,x}^{(n)}-x\right)^{2}\,\mathrm{d}t\Bigg]$
$\quad+\frac{n}{2Z_{\alpha,x}^{(n)}}\left\{\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{2}-\frac{1}{\alpha n}x(1-x)\right\}f^{\prime\prime}(x)$
$\quad+\frac{1}{2\alpha}\left(\frac{1}{Z_{\alpha,x}^{(n)}}-1\right)x(1-x)f^{\prime\prime}(x)$
$=:\mathrm{I}_{\alpha}^{(n)}(x)+\mathrm{I\hskip-1.00006ptI}_{\alpha}^{(n)}(x)+\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha}^{(n)}(x)+\mathrm{I\hskip-1.00006ptV}_{\alpha}^{(n)}(x).$ (4.5)
@@
$\quad+n\|f^{\prime}\|_{\infty}\max_{x\in[r,1-r]}\left|\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)\right|$
@@
$\quad+n\|f^{\prime}\|_{\infty}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)O(e^{-\delta n})$
@@
$\left|\mathrm{I\hskip-1.00006ptI}_{\alpha}^{(n)}(x)\right|$
@@
$\quad+\frac{n\|f^{\prime\prime}\|_{\infty}}{\rho^{2}}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{4}\right|$
@@
$\left|\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{n}{2}\left(\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\frac{1}{Z_{\alpha,x}^{(n)}}\right)\|f^{\prime\prime}\|_{\infty}O\left(\frac{1}{n^{2}}\right)=O\left(\frac{1}{n}\right)$ (4.8)
as $n\to\infty$. On the other hand, for $\varepsilon>0$, $x(1-x)<\varepsilon$ for $x\in[0,\varepsilon]\cup[1-\varepsilon,1]$. Then, we have
$\left|\mathrm{I\hskip-1.00006ptV}_{\alpha}^{(n)}(x)\right|$ $\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\max_{x\in[0,\varepsilon]\cup[1-\varepsilon,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\max_{x\in[\varepsilon,1-\varepsilon]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
$\leq\frac{\varepsilon\|f^{\prime\prime}\|_{\infty}}{2\alpha}\sup_{n\in\mathbb{N}}\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|+\frac{\|f^{\prime\prime}\|_{\infty}}{8\alpha}\sup_{\alpha\in K}\max_{x\in[\varepsilon,1-\varepsilon]}\left|\frac{1}{Z_{\alpha,x}^{(n)}}-1\right|$
@@
as $n\to\infty$, by virtue of Proposition 2.6 (1) and Theorem 2.9. Therefore, by combining (4.5) with (4.3), (4.7), (4.8), and (4.9), we obtain
@@
The function $g_{\delta}$ (resp. $\widehat{g}_{\delta}$) is non-negative and continuous on $0,\varepsilon]$ (resp. $1-\varepsilon,1]$) with $g_{\delta}(\varepsilon)=0$ (resp. $\widehat{g}_{\delta}(1-\varepsilon)=0$). We also define
@@
$G(0+)$ $=+\infty,$ $\qquad G(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(0)|,$
$\widehat{G}(0+)$ $=+\infty,$ $\qquad\widehat{G}(\varepsilon)$ $=\varepsilon\log 2-\frac{\varepsilon}{2}<|f^{\prime}(1)|$
due to (4.11). Hence, the mean value theorem implies the existence of $\delta_{0},\delta_{1}\in(0,\varepsilon)$ such that $G(\delta_{0})=|f^{\prime}(0)|$ and $\widehat{G}(\delta_{1})=|f^{\prime}(1)|$. We now define a family of continuous functions $(f_{\varepsilon})_{\varepsilon>0}\subset C([0,1])$ by
[…]
@@
[…]
@@
[…]
@@
It follows from Lemma 4.6 and [7, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$. Then, by combining this fact with Lemma 4.5, we can apply Trotter’s approximation theorem (cf. [29, Theorem 5.2]) to conclude (4.3). ∎
@@
Because (4.14) has a unique strong solution, $(X_{\alpha,x}(t))_{t\geq 0}$ is the unique solution to the martingale problem for $(\mathcal{L}_{\alpha},\delta_{x})$ for every $x\in[0,1]$. Moreover, we note that the diffusion coefficient $(2\alpha)^{-1}x(1-x)$ of $\mathcal{L}_{\alpha}$ is continuous and bounded on $[0,1]$. Then, we can apply [14, Theorem 32.11] to conclude that
@@
is a Feller semigroup on $C([0,1])$ and $\mathcal{L}_{\alpha}$ uniquely extends to the associated generator. Because a Feller semigroup is uniquely determined by its generator (cf. [14, Lemma 17.5]), we have $T_{\alpha}(t)=\widetilde{T}_{\alpha}(t)$ for all $t\geq 0$. ∎
4.4 The fractional Bernstein operator with $\alpha=\alpha(x)$ being a function of $x$
@@
$\mu_{\alpha(x),x}^{(n)}(\mathrm{d}z)=\frac{\alpha(x)}{Z_{\alpha(x),x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha(x)n}{\alpha(x)j}x^{\alpha(x)j}(1-x)^{\alpha(x)(n-j)}\delta_{j}(\mathrm{d}z)$
@@
$B_{\alpha(\cdot),n}f(x)$ $=\int_{\mathbb{R}}f\left(\frac{z}{n}\right)\,\mu_{\alpha(x),x}^{(n)}(\mathrm{d}z)$
@@
for $f\in C^{2}([0,1])$ and $x\in[0,1]$. Then, the operator $(\mathcal{L}_{\alpha(\cdot)},C^{2}([0,1]))$ is also closable on $C([0,1])$, and its closure $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$ generates the Feller semigroup $(T_{\alpha(\cdot)}(t))_{t\geq 0}$ on $C([0,1])$, which is no longer the Wright–Fisher diffusion semigroup in general (see [2, Theorems 1.6.11 and 6.2.6]).
By following the proof of Theorem 4.4, we also prove that the appropriately scaled operators $(B_{\alpha(\cdot),n}^{\lfloor nt\rfloor})_{n\in\mathbb{N}_{0}}$, $t\geq 0$, defined by the iterates of $B_{\alpha(\cdot),n}$ converges to $(T_{\alpha(\cdot)}(t))_{t\geq 0}$.
@@
as $\varepsilon\searrow 0$, which implies that $\mathcal{D}_{0}$ is also a core for $(\mathcal{L}_{\alpha(\cdot)},\mathrm{Dom}(\mathcal{L}_{\alpha(\cdot)}))$. Therefore, it follows from [7, Proposition 1.3.1] that $\mathrm{Range}(\lambda-\mathcal{L}_{\alpha(\cdot)}|_{\mathcal{D}_{0}})$ is dense in $C([0,1])$ for some $\lambda>0$. Then, Trotter’s approximation theorem allows us to obtain (4.16). ∎
@@
We assume that $\alpha=\alpha(x)\colon[0,1]\to(0,\infty)$ is continuous and $\alpha(x)^{-1}=1/\alpha(x)$ is Lipschitz continuous on $[0,1]$. For $x\in[0,1]$, let $(X_{\alpha(\cdot),x}(t))_{t\geq 0}$ be the one-dimensional diffusion process that solves the stochastic differential equation
@@
Proof.
We omit the details of the proof because it is essentially same as that of Proposition 4.7. On the other hand, we should note that (4.17) has the unique strong solution. Indeed, the existence of the weak solution to (4.17) holds true since the diffusion coefficient
$\sigma(x):=\sqrt{\alpha(x)^{-1}x(1-x)},\qquad x\in[0,1],$
of (4.17) is bounded and continuous on $[0,1]$. Moreover, it follows from the Lipschitz continuity of $1/\alpha(x)$ that
$|\sigma(x)-\sigma(y)|$ $=\left|\sqrt{\alpha(x)^{-1}x(1-x)}-\sqrt{\alpha(y)^{-1}y(1-y)}\right|$
$\leq\sqrt{\left|\alpha(x)^{-1}x(1-x)-\alpha(y)^{-1}y(1-y)\right|}$
$\leq\sqrt{\left|\alpha(x)^{-1}-\alpha(y)^{-1}\right|x(1-x)+\alpha(y)^{-1}|x(1-x)-y(1-y)|}$
$\leq\sqrt{C_{4,3}|x-y|+3\alpha_{L}^{-1}|x-y|}\leq C_{4,4}|x-y|^{1/2},\qquad x,y\in[0,1],$
which implies the pathwise uniqueness of (4.17). Therefore, the Yamada–Watanabe theorem (cf. [30, Theorem 1]) leads to the existence of the unique strong solution to (4.17). ∎
Acknowledgements. The first-named author was supported by JSPS KAKENHI Grant Numbers 19H00643 and 25K07056 and The Kyoto University Foundation. The second-named author was supported by JSPS KAKENHI Grant Numbers 23K12986 and 23K03155.
Data availability No datasets were generated or analyzed during the current study.
Declarations
Conflict of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.